OpenStudy (anonymous):

The perimeter of a rectangle is 50 cm and the length is 4 cm longer than the width. Write and solve the system of equations using the substitution method

OpenStudy (nikato):

lets call the width, w how would you write the length in terms of w

OpenStudy (texaschic101):

formula for perimeter of rectangle : P = 2L + 2W P = 50 L = W + 4 now sub 50 = 2(W + 4) + 2W now just solve for W. Once you have found W, sub that back into " L = W + 4 " and solve for L.